Angie Silva has recently opened The Sandal Shop in Brisbane, Australia, a store that specializes in fashionable sandals. In time, she hopes to open a chain of sandal shops. As a first step, she has gathered the following data for her new store: Sales price per pair of sandals $ 32 Variable expenses per pair of sandals 16 Contribution margin per pair of sandals $ 16 Fixed expenses per year: Building rental $ 9,600 Equipment depreciation 12,000 Selling 9,600 Administrative 16,800 Total fixed expenses $ 48,000 Required: 1. What is the break-even point in unit sales and dollar sales?

3. Angie has decided that she must earn a profit of $16,000 the first year to justify her time and effort. How many pairs of sandals must be sold to attain this target profit?

4. Angie now has two salespersons working in the store—one full time and one part time. It will cost her an additional $7,000 per year to convert the part-time position to a full-time position. Angie believes that the change would increase annual sales by $41,600. Should she convert the position? Use the incremental approach.

5. Refer to the original data. During the first year, the store sold only 3,500 pairs of sandals and reported the following operating results: Sales (3,500 pairs) $ 112,000 Variable expenses 56,000 Contribution margin 56,000 Fixed expenses 48,000 Net operating income $ 8,000 a. What is the store’s degree of operating leverage?

b. Angie is confident that with a more intense sales effort and with a more creative advertising program she can increase sales by 50% next year. Using the degree of operating leverage, what would be the expected percentage increase in net operating income if Angie is able to increase sales by 50%?
Answer #1

Solution:

Sales price per pair of sandals = $32

Variable expenses per pair of sandals = $16

Contribution margin = $16

Fixed Expenses per year:

Building rental = $9,600

Equipment depreciation = $12,000

Selling = $9,600

Administration = $16,800

Total Fixed expenses = $48,000

Breakeven point in unit sales can be calculated using the formula given below

Breakeven point in unit sales = Total fixed expenses / Contribution margin

                                                                  = 48000 / 16

                                                                  = 3000 units

Breakeven point in dollar sales = Breakeven point in unit sales * sales price

                                                                  = 3000 * 32

                                                                  = $96,000

3.

To earn a profit of $16,000

Sales – Total variable cost – Total fixed cost = Profit

Let the no. of sandals must be sold to attain this target profit be X

The above equation will become,

(X * Sales price per unit) – (X * variable cost per unit) – T.F.C = profit

(X *32) – (X*16) – 48000 = 16000

32 X – 16 X = 16000 + 48000

16X = 64000

X = 64000 / 16

X = 4000 units.

4000 units has to be sold to attain a profit of $16,000.

4.

If she converts the position the additional cost would be $7000 and annual sales would increase by $41,600.

No. of units’ sales increase = annual sales amount increase / sales price

                                                                 

                                                                  = 41600 / 32

                                                                  = 1300 units.

Amount going to be earned through this = 1300 * contribution margin

                                                                  = 1300* 16

                                                                  = $20,800.

She should convert the position because the amount she can make is higher than expenses incurred.

5.

a. Degree of operating leverage = (sales – variable cost) / (sales – variable cost – fixed cost)

Sales = $112,000

Variable expenses = $ 56,000

Contribution margin = $ 56,000

Fixed expenses = $ 48,000

Net Operating income = $8,000

Degree of operating leverage = (112000 – 56000) / (112000 – 56000 – 48000)

                                                                  = 56000 / 8000

Degree of operating leverage       = 7

5 b.

If sales increased by 50%, the new sales in dollar would be $168,000

New sales in unit = 168000 / 32

                               = 5250

New total variable expenses = 5250 * 16

                                                    = 84,000

Operating income = sales – T.V.E – T.F.E

                                  = 168000 – 84000 – 48000

                                  = 36000

Percentage increase in net operating income = (36000 – 8000) / 8000 *100

                                                                                       = 28000 / 8000 *100

                                                                                       = 350%